What is the pressure required to compress 2 Liters of a gas at 700 mm and 293 K into a contain of .1 capacity at a temperature of 123 K.
1
Expert's answer
2020-04-13T02:25:55-0400

In the condition of the problem the capasity of the container wasn't clearly specified. So I marked it with ' x ' liters. In order to obtaine the answer just devide the answer ( 587.71) by the right capacity.

General gas law:

p1V1/T1=p2V2/T2;

p1=700 mm Hg; V1=2 L; T1=293 K

V2=x L; T2=123 K; p2 - ?


p2=700*2*123/(x*293) = 587.71/x mm Hg